# WebQuest: A Benchmark for Multimodal QA on Web Page Sequences

### v1

@@ 4.1 Metrics
To evaluate the performance of various models on the WebQuest benchmark, we utilize a variant of the Relaxed Accuracy metric used for the ChartQA Masry et al. (2022) dataset. It is noteworthy that over half of the answers in this benchmark are numerical values. For these numerical values answers, we use heuristics to extract the numbers from predictions and ground truths. For floating-point answers, similar to Relaxed Accuracy. we allow a margin of ±0.05 in predictions to be considered correct. In contrast, for integer and string answers, we use SQuaD Rajpurkar et al. (2016) pre-processing and exact match between the prediction and ground truth as the success criterion. Furthermore, we incorporate multiple ground truth variants generated using Gemini Pro Team (2024) to accommodate various acceptable answer formats and consider the final score as the maximum score across all ground truth variants. 66todo: 6sinivasksun, jasonjlin inline,color=redinline,color=redtodo: inline,color=redsri:Did we add any evals for few shot prompting? If not, we should remove the few shot prompting mentions?
4.2 Baselines
@@
We evaluate the performance of a number of state-of-the-art MLLMs, along with the adapted forms of web-oriented models on the WebQuest benchmark. We use pretrained parameters and default settings of API-based models for evaluation. The model input includes the screen image(s) and question. Although image ordering should not be relevant to the questions, we use the same ordering of images for Multi Screen QA and the original navigation sequence orders for Trace QA. We craft the model prompts to only output only numbers for Single and Multi Screen QA and numbers or short phrases for Trace QA. For models we used out of the box, inference time techniques like prompt engineering and Chain-of-thought prompting Wei et al. (2022) were utilized to improve their performance. In our Chain-of-thought Wei et al. (2022) prompt, we include stages to analyze the screenshot and question, extract information from each screen, analyze extracted information, then finally to generate the answer. We evaluated various state of the art models like GPT-4V OpenAI et al. (2024), Gemini Flash Team (2024), Claude 3 Anthropic (2024). For Claude 3 Anthropic (2024) models, the model input is limited to 20 images. Therefore for Trace QA evaluations, we only use the last 20 images in the sequence. Furthermore, due to memory constraints for Claude 3 Anthropic (2024) models, we resize the images preserving aspect ratio such that the shorter side is 1440 pixels for Trace QA, whereas we use the original image resolutions for Single screen QA and Multi screen QA. For GPT-4V OpenAI et al. (2024) and Gemini Flash Team (2024) evals we use the original image resolutions for all questions.
@@
Claude-3 Sonnet 27.3 38.9 13.9 41.9 23.2 19.6
Claude-3 Opus 25.3 40.4 15.1 23.3 18.5 26.6
Gemini Flash 34.9 48.9 11.1 46.8 24.6 35.9
GPT4-V 36.7 57.4 24.1 46.5 29.0 40.8
Table 1: Performance of various MLLMs on variants of the WebQuest benchmark. All these models are used in a zero shot setting without any fine-tuning. For each of the models, prompt tuning was performed separately. Overall, we can see that all of the MLLMs perform a lot better on the Single Screen QA split and their performance decreases as the number of screens increases. Also, note that the performance of the MLLMs on the Trace QA split is considerably lower than on the Multi Screen QA split.
Figure 8: Performance of different models vs number of screens in each question. Perhaps surprisingly, we do not see clear evidence of model performance reducing as number of screens increases.
Figure 9: Performance of different models for different question types. The types ending in a + sign e.g., arithmetic+ indicate questions which contain multiple types with arithmetic being one of them. Here we notice that the different models do not perform well on questions involving counting, whereas they do well on arithmetic questions. All models presented here were evaluated with CoT prompts.
4.3 Main Results
@@
4.3.1 Single Screen QA
As described in section 3, the single screen question answering split involves answering a question based on the content of a single web screen. On this task, we notice that all the MLLMs perform relatively well compared to Multi Screen QA and Trace QA. Chain-of-thought Wei et al. (2022) prompting improves the scores significantly. We also note that the performance difference for large MLLMs is very close on this split when compared to a model like ScreenAI Baechler et al. (2024) specifically trained on UI data.
4.3.2 Multi Screen QA
To recap briefly, the multi screen question answering split involves question answering across various relevant screens from different websites. As we saw in section 3, the images are from related websites and on average, each QA pair contains 3.84 screens with 3, 4 or 5 screens. For this split, we notice that all the MLLMs perform worse than the single screen split. With some MLLMs, we notice considerable performance improvement with Chain-of-thought Wei et al. (2022) prompting and few shot evaluations. However, the performance for all the MLLMs drops to around 45% on this split, highlighting the difficulty some of these models have in reasoning across multiple images. We further analyze the performance gap in section 4.4.
4.3.3 Trace QA
The main difference between this split and the Multi Screen QA 4.3.2 is that this split also contains the screens seen during navigation to the relevant screens for answering the questions. As stated in section 3, the Multi Screen QA split is a simplified version of this split as it already contains only the relevant screens. This split can also be used for evaluating different agent systems since it contains action information for each step enabling step-by-step supervision compared to existing datasets. On average, each question has $15.815.8$ 77todo: 7Update this number from the stats screens, out of which relevant information is present on $3.23.2$ color=redcolor=redtodo: color=redUpdate this number from the stats - how do we compute this? screens.
We notice that the performance of all the models is significantly worse on this split compared to both the Single Screen and Multi Screen QA splits. With Chain-of-thought prompting Wei et al. (2022) and prompt engineering, MLLMs on average achieve a score of 30.6%. Out of the MLLMs we evaluated, GPT-4V OpenAI et al. (2024) achieves the best performance with a score of 40.8% on this dataset. Claude 3 Anthropic (2024) models performs significantly worse than GPT-4V OpenAI et al. (2024) and Gemini Flash Team (2024) models, likely due to the extra image resizing step to avoid out of memory errors.
4.4 Analysis
@@
We observe that all the models including the largest ones achieve much higher performance on single screen tasks compared to multi screen tasks. However, as we can see in figure 8, for the MultiScreen QA split, there is no clear evidence of model performance decreasing as the number of relevant screens increase. This observation needs further investigation. From figure 9 we observe a significant performance gap in questions involving counting across screens compared to arithmetic or comparison operations.
Impact of prompting techniques on results
@@
We observe that prompting techniques like Chain-of-thought Wei et al. (2022), asking the models to split the question into its constituent parts and describing the contents of each screen results in significant improvements across different kinds of models, even with the largest models across the GPT-4 OpenAI et al. (2024), Gemini Team (2024) and Claude 3  Anthropic (2024) model families. On average, we see improvements of 26.5 percentage points on Multi screen QA split and 6.7 percentage points on the Trace QA split. When we manually observe the CoT responses generated, we do note that the model is able to answer sub-questions on various screens correctly and this ability results in better overall performance.
Varying performance on different question categories
@@
For each of the questions in the dataset, we also provide information on the kind of reasoning needed to answer the question. Inspired by similar annotations in the InfographicVQA Mathew et al. (2022) dataset, we use the categories: arithmetic, comparison, counting and filtering where each question can have more than one category. As shown in figure 9, we note that different models exhibit different performances across the various categories. Across all models, questions related to counting tend to get a lower accuracy than other ones, while questions involving only arithmetic operations get a higher score. However, it is important to note that these aggregate numbers relying on the final answer are conflating errors due to reasoning with errors due to information extraction or other causes. We also note that the performance decreases as the number of entities across the screens increases indicating the model may not able to extract all the information correctly when a large number of entities are present.
5 Conclusion
@@
In this work, we introduce WebQuest, the first multimodal question answering benchmark designed to evaluate reasoning abilities across multiple screens. Our benchmark encompasses three distinct settings: single-screen, multi-screen, and trace-based, enabling a comprehensive assessment of model capabilities. Evaluations of state-of-the-art MLLMs, including Gemini Flash Team (2024), Claude 3 Anthropic (2024), and GPT-4V OpenAI et al. (2024), reveal a significant performance gap between single-screen and multi-screen reasoning tasks. Notably, Chain-of-thought prompting Wei et al. (2022) proves effective for navigating complex, multi-screen scenarios by facilitating information extraction and synthesis across multiple screens. Future research directions include leveraging richer screen information, such as OCR, DOM, and screen annotations from tools like ScreenAI Baechler et al. (2024), to potentially enhance model performance and even replace raw screen images in cases of long screen sequences. Furthermore, extending WebQuest to support personalized, multi-turn dialogue scenarios presents a compelling avenue for future exploration, enabling the development of agents capable of assisting users in dynamic, context-aware conversations.
@@ Appendix A Dataset Examples
We show a few examples of the WebQuest dataset, in Figure 10 for Single Screen QA, in Figures 11 and 12 for Multi Screen QA, and in Figure 13 for Trace QA.
@@
Figure 10: Examples of Single Screen QA.
@@
Figure 11: An example of Multipage QA, about prices of selected gym accessories, each from a different online shop. Notice the layouts and font sizes vary among them.
@@
Figure 12: An example of Multipage QA, about prices of food items over 4 websites of restaurant menus. Notice the layouts and font sizes vary significantly among them.
@@
Figure 13: An example of Trace QA. The example, situated in a rental exploration scenario, is about price differences between different options. Notice the trace captures the steps taken in the exploration.
@@ Appendix B Statistics of the Dataset
We show detailed statistics of the dataset in Table 4, Table 4, and Table 4.
Category Num Examples
Restaurant 44
Shopping 184
Hotel 68
Transportation 114
Landmark 21
News 22
Media Sharing 7
Education & Information 9
Movies 24
Real Estate 19
Other 30
Total 542
Table 2: Distribution of the 542 examples of Single Screen QA, per categories and the number of screens in the example.
3 screens 4 screens 5 screens
54 90 27
Table 3: Distribution of the 171 examples of Multi Screen QA, per categories and the number of screens in the example.
@@
Table 4: Distribution of examples of Trace QA for each category

### v2

@@ 4.1. Metrics
To evaluate the performance of various models on the WebQuest benchmark, we utilize a variant of the Relaxed Accuracy metric used for the ChartQA (Masry et al., 2022) dataset. It is noteworthy that over half of the answers in this benchmark are numerical values. For these numerical value answers, we use heuristics to extract the numbers from predictions and ground truths. For floating-point answers, similar to Relaxed Accuracy, we allow a margin of ±0.05 in predictions to be considered correct. In contrast, for integer and string answers, we use SQuaD (Rajpurkar et al., 2016) pre-processing and exact match between the prediction and ground truth as the success criterion. Furthermore, we incorporate multiple ground truth variants generated using Gemini Pro (Team, 2024) to accommodate various acceptable answer formats and consider the final score as the maximum score across all ground truth variants.
4.2. Baselines
@@
We evaluate the performance of a number of state-of-the-art MLLMs and leading open source models on our benchmark. We use pretrained parameters and default settings of API-based models for evaluation. The model input includes the screen image(s) and question. Although image ordering should not be relevant to the questions, we use the same ordering of images for Multi Screen QA and the original navigation sequence orders for Trace QA. For models we used out of the box, inference time techniques like prompt engineering and Chain-of-thought prompting (Wei et al., 2022) were utilized to improve their performance. In our Chain-of-thought (Wei et al., 2022) prompt, we include stages (steps) to analyze the screenshot and question, extract information from each screen, analyze extracted information, then finally to generate the answer in short phrases. We evaluated various state of the art models like GPT-4V (OpenAI et al., 2024), Gemini Flash (Team, 2024), Claude 3 (Anthropic, 2024). For Claude 3 (Anthropic, 2024) models, the model input is limited to 20 images. Therefore for Trace QA evaluations, we only use the last 20 images in the sequence. Furthermore, due to memory constraints for Claude 3 (Anthropic, 2024) models, we resize the images preserving aspect ratio such that the shorter side is 1440 pixels for Trace QA, whereas we use the original image resolutions for Single Screen QA and Multi Screen QA. For GPT-4V (OpenAI et al., 2024) and Gemini Flash (Team, 2024) evals we use the original image resolutions for all questions. We also evaluate open sourced models, including BLIP2 (Li et al., 2023), InstructBLIP (Dai et al., 2023) and PaLIGemma (Beyer et al., 2024). We evaluate BLIP2 and InstructBLIP in a zero-shot setting, and in both cases, use the variant based on Flan-T5-XXL language model (Chung et al., 2022). Following the recommendations from the official documentation (for Developers, 2024), we evaluate a fine-tuned version of PaliGemma. We fine-tune the pre-trained pt-896 checkpoint on a mixture of open source datasets: DocVQA (Mathew et al., 2021), mobile ScreenQA (Hsiao et al., 2022), ChartQA (Masry et al., 2022) and InfographicVQA (Mathew et al., 2022). We set 1e-5 as the initial learning rate, with a warmup of 0.05% and a cosine decay. We run fine-tuning for 10 epochs.
Table 1. Performance of various MLLMs on variants of the WebQuest benchmark. The metric used here is Relaxed Accuracy. All these models, except PaLIGemma are used in a zero shot setting without any fine-tuning. For each of the models, prompt tuning was performed separately. Overall, we can see that all of the MLLMs perform a lot better on the Single Screen QA split and their performance decreases as the number of screens increases. Also, note that the performance of the MLLMs on the Trace QA split is considerably lower than on the Multi Screen QA split.
@@
Proprietary models
Claude-3 Sonnet 27.3 38.9 15.4 31.2 23.2 19.6
Claude-3 Opus 25.3 40.4 16.3 23.4 18.5 26.6
Gemini Flash 34.9 48.9 15.8 31.1 24.6 35.9
GPT-4V 36.7 57.4 23.2 44.0 29.0 40.8
Open-source models
PaLIGemma 10.5 N/A <5% N/A N/A N/A
BLIP-2 6.7 N/A N/A N/A N/A N/A
InstructBLIP 9.1 N/A 8.8 N/A N/A N/A
Figure 10. Performance of different models vs number of screens in each question. Perhaps surprisingly, we do not see clear evidence of model performance reducing as number of screens increases.
Figure 11. Performance of different models for different question types. The types ending in a + sign e.g., arithmetic+ indicate questions which contain multiple types with arithmetic being one of them. Here we notice that the different models do not perform well on questions involving counting, whereas they do well on arithmetic questions. All models presented here were evaluated with CoT prompts.
4.3. Main Results
@@
4.3.1. Single Screen QA
As described in section 3, the single screen question answering split involves answering a question based on the content of a single web screen. On this task, we notice that all the MLLMs perform relatively well compared to Multi Screen QA and Trace QA. Chain-of-thought (Wei et al., 2022) prompting improves the scores on average 15.4%. We also notice a significant performance difference between proprietary models and smaller open source models PaLIGemma and InstructBLIP.
4.3.2. Multi Screen QA
To recap briefly, the multi screen question answering split involves question answering across various relevant screens from different websites. As we saw in section 3, the images are from related websites and on average, each QA pair contains 4.22 screens with 3, 4 or 5 screens. For this split, we notice that all the MLLMs perform worse than the single screen split. With some MLLMs, we notice considerable performance improvement with Chain-of-thought (Wei et al., 2022) prompting and few shot evaluations. However, the performance for all the MLLMs drops to on average 32.4% on this split, highlighting the difficulty some of these models have in reasoning across multiple images. We further analyze the performance gap in section 4.4. For open-source models, similar to Single Screen QA, we notice a significant performance gap to proprietary models. Also, we could not adapt BLIP2 model to run inference with multiple images, so we do not report any numbers for this split. As for PaLIGemma, since it was fine-tuned on single page VQA datasets (including DocVQA (Mathew et al., 2021), mobile ScreenQA (Hsiao et al., 2022), ChartQA (Masry et al., 2022) and InfographicVQA (Mathew et al., 2022)), the performance is very low.
4.3.3. Trace QA
The main difference between this split and the Multi Screen QA 4.3.2 is that this split also contains the screens seen during navigation to the relevant screens for answering the questions. As stated in section 3, the Multi Screen QA split is a simplified version of this split as it already contains only the relevant screens. This split can also be used for evaluating different agent systems since it contains action information for each step enabling step-by-step supervision compared to existing datasets. On average, each question has $15.815.8$ screens, out of which relevant information is present on $3.23.2$ screens. We notice that the average performance of all the models is worse on this split compared to both the Single Screen and Multi Screen QA splits. With Chain-of-thought prompting (Wei et al., 2022) and prompt engineering, MLLMs on average achieve a score of 30.6%. Out of the MLLMs we evaluated, GPT-4V (OpenAI et al., 2024) achieves the best performance with a score of 40.8% on this dataset. Claude 3 (Anthropic, 2024) models performs significantly worse than GPT-4V (OpenAI et al., 2024) and Gemini Flash (Team, 2024) models, likely due to the extra image resizing step to avoid out of memory errors. For open-source models, the large number of input images per example resulted in out of memory errors.
4.4. Analysis
@@
We observe that all the models including the largest ones achieve much higher performance on single screen tasks compared to multi screen tasks. However, as we can see in figure 11, for the MultiScreen QA split, there is no clear evidence of model performance decreasing as the number of relevant screens increase. From figure 11 we observe a significant performance gap in questions involving counting across screens compared to arithmetic or comparison operations.
4.4.1. The impact of Chain-of-thought prompting
@@
We observe that Chain-of-thought (Wei et al., 2022) prompting results in significant improvements across all models, even with the largest models across the GPT-4 (OpenAI et al., 2024), Gemini (Team, 2024) and Claude 3 (Anthropic, 2024) model families. On average, we see improvements of 15% on Multi screen QA split and 6.7% on the Trace QA split. When we manually inspect model step-by-step responses, we observe that the model is able to answer sub-questions on various screens correctly and this ability results in better overall performance. We also observed Chain-of-thought prompting (Wei et al., 2022) improves multi-page to a larger extent than single-screen tasks (avg. 20.3% vs 15.4%) for most models. Claude-3 Opus (Anthropic, 2024), which had the second highest baseline performance without Chain-of-thought prompting, ranked last among proprietary models when evaluated with CoT. One of the causes of reduced performance was that, with Chain-of-thought prompting, Claude-3 Opus produced N/A 11% of the time in Multi-page QA.
4.4.2. Model performance on Chain-of-thought reasoning steps
In our Chain-of-thought (Wei et al., 2022) prompt, we split the original questions to 4 reasoning steps, including question analysis, screenshot analysis, information extraction, analyze information, and answer generation. We observe that all proprietary models perform well for question analysis, as illustrated in  15. For screenshot analysis, some models focus on higher-level semantics with common sense context while successful examples showed more attention to detail. i.e. "The second screenshot is a website for a restaurant called Vic’s, showing popular menu items and their prices." vs. "Vic’s restaurant with items like Cacio e Pepe for \$20.00, Roasted Chicken for \$28.00, and Rainbow Cookie Plate for \$8.00". Despite both models being successful in information extraction, the model that is more abstract in screenshot analysis performed the incorrect arithmetic and produced the wrong answer, and the correct model’s screenshot analysis were more direct. Interestingly, we found some models to consistently do better at one and worse at the other step, but we observed no strong correlation between final answer correctness with any specific steps.
The wine example in Figure 13 illustrates some interesting differences between even the same family of models. Claude-3 Sonnet (Anthropic, 2024) does not recognize the food in page 1 while Claude-3 Opus (Anthropic, 2024) and Gemini Flash does, suggesting they leverage strong common sense knowledge. However, all models fail to recognize the color and font difference of the menu item subheadings, which are attributes of the drink selections and should not be counted. Gemini Flash was poorer at counting, while understanding the task better. GPT-4V (OpenAI et al., 2024) tends to hallucinate and make the incorrect associations in its reasoning attributed to reading error (logical fallacy), i.e. mistaking price as number of options and missing the wine option on page 3.
Question: The prices of lasagna and Carolina BBQ Oink Sampler together are how much more than the prices of 3 tenders and 2 pieces of fish together?
Answer: \$493.02
Figure 12. An example where UI differences, despite serving the same checkout functionality of a shopping cart, can confound models even with Chain-of-Thought prompting.
Question: Total, how many drinks are shown on these site pages?
Answer: 18
Figure 13. A challenging example that all models failed to answer correctly. The counting task is complicated by the stylistic diversity of the menu. Each models’ Chain-of-thought gets it wrong at different steps.
Figure 14. Models fails to list correctly the dates with the lowest temperature. This involves first associating each date with its corresponding temperature and then going through each one to decide which one matches the criteria.
Figure 15. With Chain-of-thought prompting, Gemini Flash extracts relevant information from the screen sequence, then generates the correct answer.
4.4.3. Common errors made by proprietary models
  • Analysis of UI Interfaces We noticed all proprietary models tend to be confused by visual differences in UI interfaces serving the same purpose. As displayed in Figure  13, when presented with multiple checkout screens, one displaying itemized prices and the other one only the total price, even when GPT-4V misread the individualized price as total price, an example of visual oversight.
  • Counting instances under a given criteria Another common failure mode is counting instances under a given criteria or association. These questions often involve counting and filtering based on both semantic and visual attributes, such as dining options, costs, colors and style. (See Figure 14 for an example.) We found Chain-of-thought step hints to be particularly helpful on decoupling conflated errors across information extraction and reasoning, providing tremendous value in grounding the model’s errors and comparing performance between models. An example is illustrated at Figure  13.
  • Considering questions non answerable We also observe many errors to be "the question is not answerable" across proprietary models. In these cases, we observe that the main reason is that the model fails to identify relevant information from the screens
4.4.4. Varying performance on different question categories
@@
For each of the questions in the dataset, we also provide information on the kind of reasoning needed to answer the question. Inspired by similar annotations in the InfographicVQA (Mathew et al., 2022) dataset, we use the categories: arithmetic, comparison, counting and filtering where each question can have more than one category. We use Gemini 1.5 Pro (Team et al., 2023) to optionally assign one or more of the categories to each question. We manually inspect all the categories assigned and make necessary corrections. As shown in figure 11, we note that different models exhibit different performances across the various categories. Across all models, questions related to counting tend to get a lower accuracy than other ones, while questions involving only arithmetic operations get a higher score. However, it is important to note that these aggregate numbers relying on the final answer are conflating errors due to reasoning with errors due to information extraction or other causes. We also note that the performance decreases as the number of entities across the screens increases indicating the model may not able to extract all the information correctly when a large number of entities are present.
4.4.5. Open source model performance
We observe that, open source models perform worse than proprietary LLMs on WebQuest tasks. We believe that the difference in the sizes of the open source models compared to proprietary LLMs is a significant factor here. We observed that for BLIP-2 variants  (Li et al., 2023; Dai et al., 2023), Chain-of-thought prompting (Wei et al., 2022) did not result in a more detailed output and hence show no performance improvement compared to prompts with no CoT. We also note that InstructBLIP (Dai et al., 2023) was trained with an image resolution of 224x224. However, WebQuest has screenshots with significantly higher resolutions and requires the understanding of the fine grained details of text and UI elements present on the screen. Unsuprisingly, we noticed performance improvement with inference at a resolution of 448x448 but did not see further improvements with increasing image resolution.
5. Conclusion
@@
In this work, we introduce WebQuest, the first multimodal question answering benchmark designed to evaluate reasoning abilities across multiple screens. Our benchmark encompasses three distinct settings: single-screen, multi-screen, and trace-based, enabling a comprehensive assessment of model capabilities. Evaluations of state-of-the-art MLLMs, including Gemini Flash (Team, 2024), Claude 3 (Anthropic, 2024), and GPT-4V (OpenAI et al., 2024), reveal a significant performance gap between single-screen and multi-screen reasoning tasks. Notably, Chain-of-thought prompting (Wei et al., 2022) proves effective for navigating complex, multi-screen scenarios by facilitating information extraction and synthesis across multiple screens. Future research directions include leveraging richer screen information, such as OCR, DOM, and screen annotations from tools like ScreenAI (Baechler et al., 2024), to potentially enhance model performance and even replace raw screen images in cases of long screen sequences. Furthermore, extending WebQuest to support personalized, multi-turn dialogue scenarios presents a compelling avenue for future exploration, enabling the development of agents capable of assisting users in dynamic, context-aware conversations. 11todo: 1sri: Maybe add a line about memory or retrieval. Mention this as a potential future research direction.
Limitations and Risks
Some of the limitations of our work which we plan to address in future research are:
  • Our current dataset does not include multi-turn interactions which contain follow up questions or clarifications to previous questions.
  • To avoid any personally identifiable information in the screen data, we do not include screens that require user specific credentials.
  • The question and answers, navigation traces collected may reflect cultural and demographic biases of the raters.
APPENDIX
@@ Appendix A Dataset Examples
We show a few examples of the WebQuest dataset, in Figure 16 for Single Screen QA, in Figures 17 and 18 for Multi Screen QA, and in Figure 19 for Trace QA.
@@
Question: What is the maximum discount?
Answer: 15% off
(c) The task is retrieve all the discounts on the screen and compare them against each other.
Figure 16. Examples of Single Screen QA.
@@
Figure 17. An example of Multipage QA, about prices of selected gym accessories, each from a different online shop. Notice the layouts and font sizes vary among them.
@@
Figure 18. An example of Multipage QA, about prices of food items over 4 websites of restaurant menus. Notice the layouts and font sizes vary significantly among them.
@@
Figure 19. An example of Trace QA. The example, situated in a rental exploration scenario, is about price differences between different options. Notice the trace captures the steps taken in the exploration.
@@ Appendix B Statistics of the Dataset
We show detailed statistics of the dataset in Table 5, Table 5, Table 5, and Table 5.
Category Num Examples Restaurant 44 Shopping 184 Hotel 68 Transportation 114 Landmark 21 News 22 Media Sharing 7 Education & Information 9 Movies 24 Real Estate 19 Other 30 Total 542 Table 2. Distribution of the 542 examples of Single Screen QA, per categories and the number of screens in the example. Category Num Examples Shopping 209 Restaurant 72 Transportation 12 Hotel 12 Landmark 1 Other 1 Total 307 Table 3. Distribution of the 307 examples of Multi Screen QA per categories.
3 screens 4 screens 5 screens Total
54 131 122 307
Table 4. Distribution of the 307 examples of Multi Screen QA per the number of screens in the example.
@@
Table 5. Distribution of examples of Trace QA for each category